A population consists of the five measurements 2, 6, 8, 0, and 1.



1. How many different samples of size 3 can be drawn from the population (no replacement)?

2. Construct the sampling distribution of the sample means.


1
Expert's answer
2021-06-03T05:04:39-0400

 1.

We have population values "2,6,8,0,1" population size "N=5" and sample size "n=3."Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{3}=10"


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,6, 8 & 16/3 \\
 \hdashline
 2 & 2,6,0 & 8/3 \\
 \hdashline
 3 & 2,6,1 & 9/3 \\
 \hdashline
 4 & 2,8,0 & 10/3\\
 \hdashline
 5 & 2,8,1 & 11/3 \\
\hdashline
 6 & 2,1,0 & 3/3 \\
 \hline
7 & 6,8,1 & 15/3 \\
 \hline
 8 & 6,8,0 & 14/3 \\
 \hline
9 & 6,1,0 & 7/3 \\
 \hline
10 & 8,1,0 & 9/3 \\
 \hline
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 3/3 & 1& 1/10 & 3/30 & 9/90 \\
 \hdashline
 7/3 & 1 & 1/10 & 7/30 & 49/90 \\
 \hdashline
 8/3 & 1 & 1/10 & 8/30 & 64/90 \\
 \hdashline
 9/3 & 2 & 2/10 & 18/30 & 162/90 \\
 \hdashline
 10/3 & 1 & 1/10 & 10/30 & 100/90 \\
 \hdashline
 11/3 & 1& 1/10 & 11/30 & 121/90 \\
 \hdashline
 14/3 & 1& 1/10 & 14/30 & 196/90 \\
 \hdashline
 15/3 & 1& 1/10 & 15/30 & 225/90 \\
 \hdashline
 16/3 & 1& 1/10 & 16/30 & 256/90 \\
 \hdashline
 Total & 10 & 1 & 102/30 & 1182/90 \\ \hline
\end{array}"



3.

Mean


"\mu=\dfrac{2+6+8+0+1}{5}=3.4"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{102}{30}=3.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.4=\mu"


4.

Variance


"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-2.4)^2+(8-3.4)^2""(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.0725"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{1182}{90}-(\dfrac{102}{30})^2=\dfrac{118}{75}\approx1.573333"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{118}{75}}\approx1.254326"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{9.44}{3}(\dfrac{5-3}{5-1})""=\dfrac{9.44}{6}=\dfrac{118}{75}\approx1.573333, True"